# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 6 The Auxiliary dpwpa for the GTESs $E$ and $F$
Let $E$ and $F$ be GTESs over a signature $\Sigma$. $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$ if and only if there is a path of positive length or a cycle from the vertex $a$ to the vertex $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
Proof.
By Lemma 6.66 and Lemma 6.69, we have the lemma. ∎
@@ 7 Running Examples Further Continued
We now continue Examples 3.17 and 5.50. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F]$ for the GTESs $E$ and $F$. Recall that
@@
Observe that $A[E;F]=\{\,(f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
@@
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, and
    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$.
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$.
@@
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$, and
    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$.
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$, and
    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
@@
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$   and   #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
@@
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$   and   #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
@@
We now continue Example 3.21 and its sequel 5.54. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. By definition,
$C\langle E\cup F,\emptyset\rangle,A[E;F]\rangle=C\langle E\cup F,\emptyset\rangle$.
@@
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle E,F\rangle$ hence #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$,
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle F,E\rangle$ hence #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
By direct inspection of $C\langle E,F\rangle$, $R\langle E,F\rangle$, $C\langle E\cup F,\emptyset\rangle$, and $R\langle E\cup F,\emptyset\rangle$, we get that $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, hence #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
The rewrite rule $f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ is in $R\langle E\cup F,\emptyset\rangle$ and $f(\#,\$)/_{\Theta\langle F,\emptyset\rangle}\subseteq\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. However, in $R\langle F,E\rangle$ there does not exist a rewrite rule with left-hand side $f(f(\#,\$)/_{\Theta\langle F,\emptyset\rangle},f(\#,\$)/_{\Theta\langle F,\emptyset\rangle})$. Hence $R\langle F,E\rangle$ does not keep up with $R\langle E\cup F,\emptyset\rangle$ writing #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
@@
We now continue Examples 3.22 and 5.55. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Observe that
$A[E;F]=\{\,(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\$/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\$/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
@@
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$,
We now show that GTRS $R\langle F,\emptyset\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
  • #$\#/_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Furthermore, #$\#/_{\Theta\langle F,E\rangle}$ is the only element of $C\langle F,E\rangle$ which is a subset of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
  • For the rewrite rule #˙#$\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$, there is a rewrite rule with left-hand side ##\## in $R\langle F,E\rangle$.
@@
Consequently, GTRS $R\langle F,E\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing ##\##, and
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
We now show that GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
  • $g(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$,
  • $g(g(\$))/_{\Theta\langle F,E\rangle})\subseteq\$/_{\Theta\langle E\cup F,\emptyset\rangle})$, and
  • in $R\langle E\cup F,\emptyset\rangle$, there does not exist a rewrite rule with left-hand side $g(g(g(\$))/_{\Theta\langle F,E\rangle})$.
Thus GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$, and
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
@@ 8 Constructing the Auxiliary dpwpa for the GTESs $E$ and $F$
We construct the sets and relations and dpwpa, which we introduced in Section 6 and our decision procedure needs.
[…]
@@
  • $INC(E\cup F,E)$ stored without duplicates in a red-black tree saved in a variable $RBT1$, for every ordered pair$(a,b)\in INC\langle E\cup F,E\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
  • $INC(E\cup F,F)$ stored in a red-black tree saved without duplicates in a variable $RBT2$, for every ordered pair$(a,b)\in INC(E\cup F,F)$, the search key is the first component $a$, the second component, $b$ is satellite data;
  • $NUM(E\cup F,E)$ stored in a red-black tree saved without duplicates in a variable $RBT3$, for every ordered pair$(a,k)\in NUM(E\cup F,E)$, the search key is the first component $a$, the second component, $k$ is satellite data;
  • $NUM(E\cup F,F)$ stored in a red-black tree saved without duplicates in a variable $RBT4$ for every ordered pair$(a,k)\in NUM(E\cup F,F)$, the search key is the first component $a$, the second component, $k$ is satellite data;
@@
  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved without duplicates in a variable $RBT6$; for every ordered pair$(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
@@
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$;
@@
$a,a_{1},\ldots,a_{mar_{\Sigma}}:C\langle E\cup F,\emptyset\rangle$;  $b,b_{1},\ldots,b_{mar_{\Sigma}}:C\langle E,F\rangle$;   $c,c_{1},\ldots,c_{mar_{\Sigma}}:C\langle F,E\rangle$;
@@
1. Create a subterm dag $G=(V,A,\lambda)$ for $E$ and $F$.
@@
By Proposition 2.16, compute the congruence closure $\rho\langle F,E\rangle$ of the relation $\tau\langle F,E\rangle=\{\,(u,v)\mid(\hat{u},\hat{v})\in W\,\}.$
@@
for each $c\in C\langle E\cup F,\emptyset\rangle$ do
@@
$c.equal_{E}=\mathrm{\mathrm{false}};$;
$c.equal_{F}=\mathrm{\mathrm{false}};$
@@
$a:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
$b:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x)$;
$c:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x)$;
$INC(E\cup F,E):=INC(E\cup F,E)\cup\{\,(c,a)\,\}$;
{ store $INC(E\cup F,E)$ without duplicates in the red-black tree $RBT1$.}
$INC(E\cup F,F):=INC(E\cup F,F(\cup\{\,(c,b)\,\}$;
{ store $INC(E\cup F,F)$ without duplicates in the red-black tree $RBT2$.}
if $|a|=|c|$
${\bf then}$ $c.equal_{E}:=\mathrm{\mathrm{true}};$
if $|b|=|c|$
${\bf then}$ $c.equal_{F}:=\mathrm{\mathrm{true}};$
{ By Proposition 2.16 we know the cardinalities of all $W/_{\Theta\langle E,F\rangle}$-classes, $W/_{\Theta\langle F,E\rangle}$-classes, and $W/_{\Theta\langle E\cup F,\emptyset\rangle}$-classes. }
@@
$a_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
$b_{i}:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x/i)$;
$c_{i}:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x/i)$;
@@
if $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\not\in R\langle E\cup F,\emptyset\rangle$
@@
$R\langle E\cup F,\emptyset\rangle:=R\langle E\cup F,\emptyset\rangle\cup\{\,\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\,\}$;
{ Store $R\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT5$. }
$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}:=0$;
$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{F}:=0$;
$BSTEP\langle E\cup F,\emptyset\rangle:=BSTEP\langle E\cup F,\emptyset\rangle\cup\{\,(c_{1},c),\ldots,(c_{m},c)\,\}$;
{ Store $BSTEP\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT6$}
@@
if $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\not\in R\langle E,F\rangle$
@@
$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\,\}$;
{ Store $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT7$. }
$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}:=\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}+1$;

### v2

@@ 6 The Auxiliary dpwpa for the GTESs $E$ and $F$
Let $E$ and $F$ be GTESs over a signature $\Sigma$. $R\langle E\cup F,\emptyset\rangle$ reaches a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$ if and only if there is a path of positive length or a cycle from the vertex b{b} to the vertex c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
@@ 7 Running Examples Further Continued
Observe that $A[E;F]=\{\,(f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
@@
    $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,  $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, and
    $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$.
    $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,  $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
    $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$.
@@
    $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,  $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$, and
    $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$.
    $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,  $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$, and
    $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
@@
We now continue Examples 3.19 and 5.52. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Recall that
@@
    $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$   and   $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
@@
    $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$   and   $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
@@
We now continue Example 3.22 and its sequel 5.55. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Recall that
[…]
@@
$_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle E,F\rangle$ hence $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$,
$_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle F,E\rangle$ hence $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
By direct inspection of $C\langle E,F\rangle$, $R\langle E,F\rangle$, $C\langle E\cup F,\emptyset\rangle$, and $R\langle E\cup F,\emptyset\rangle$, we get that $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $_{\Theta\langle E\cup F,\emptyset\rangle}$, hence $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
The rewrite rule $f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ is in $R\langle E\cup F,\emptyset\rangle$ and $f(\#,\mathdollar)/_{\Theta\langle F,E\rangle}\subseteq\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. However, in $R\langle F,E\rangle$ there does not exist a rewrite rule with left-hand side $f(f(\#,\mathdollar)/_{\Theta\langle F,E\rangle},f(\#,\mathdollar)/_{\Theta\langle F,E\rangle})$. Hence $R\langle F,E\rangle$ does not keep up with $R\langle E\cup F,\emptyset\rangle$ writing $_{\Theta\langle E\cup F,\emptyset\rangle}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
@@
We now continue Examples 3.23 and 5.56. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Observe that
$A[E;F]=\{\,(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
@@
$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,
$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$,
We now show that GTRS $R\langle F,E\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
  • 1.
    $_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Furthermore, $_{\Theta\langle F,E\rangle}$ is the only element of $C\langle F,E\rangle$ which is a subset of $_{\Theta\langle E\cup F,\emptyset\rangle}$.
  • 2.
    For the rewrite rule $\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$, there is a rewrite rule with left-hand side #\# in $R\langle F,E\rangle$.
  • 3.
@@
Consequently, GTRS $R\langle F,E\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing #\#, and
$_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
We now show that GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
  • 1.
    $g(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$,
  • 2.
    $g(g(\mathdollar))/_{\Theta\langle F,E\rangle})\subseteq\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})$, and
  • 3.
    in $R\langle E\cup F,\emptyset\rangle$, there does not exist a rewrite rule with left-hand side $g(g(g(\mathdollar))/_{\Theta\langle F,E\rangle})$.
Thus GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$, and
$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
@@ 8 Constructing the Auxiliary dpwpa for the GTESs $E$ and $F$
We adopt and slightly modify the fast ground completion algorithm of Fülöp and Vágvölgyi in Definition 4.29. This way, we produce the GTRSs $R\langle E,F\rangle$, GTRS $R\langle F,E\rangle$, and $R\langle E\cup F,\emptyset\rangle$, and construct the sets and relations and dpwpa, which we introduced in Section 6 and our decision procedure needs.
We define a mapping $counter_{E}:R\langle E\cup F,\emptyset\rangle\rightarrow{\mathbb{N}}$, to each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$, $m\in{\mathbb{N}}$, in $R\langle E\cup F,\emptyset\rangle$, we assign the number of rules $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ in $R\langle E,F\rangle$, where ${c}_{i}\subseteq{b}_{i}$ for every $i\in\{\,1,\ldots,m\,\}$. In pseudocode we treat the nonnegative integer $counter_{E}(\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b})$ as satellite data of the rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$, and denote it as $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}.counter_{E}$. In this way, we store the mapping $counter_{E}$. We define the mapping $counter_{F}:R\langle E\cup F,\emptyset\rangle\rightarrow{\mathbb{N}}$ and treat $counter_{F}(\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b})$ in a similar way.
@@
  3. 3.
[…]
  4. 4.
[…]
  5. 5.
[…]
  6. 6.
[…]
  7. 7.
@@
  8. 8.
[…]
  9. 9.
@@
  11. 11.
    the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$;
@@
${b},{b}_{1},\ldots,{b}_{mar_{\Sigma}}:C\langle E\cup F,\emptyset\rangle$;  ${c},{c}_{1},\ldots,{c}_{mar_{\Sigma}}:C\langle E,F\rangle$;   ${d},{d}_{1},\ldots,{d}_{mar_{\Sigma}}:C\langle F,E\rangle$;
@@
(Here comes the main part of the algorithm CAD. We run the fast ground completion algorithm in Definition 4.29 for each of GTESs $E$, $F$, and $E\cup F$, in this way we compute data items 1, 2, 7, 9, and 10 above. During the run of the algorithm CAD, it computes also the other data items listed above. Our algorithm consists of six steps:
  1. 1.
    We create a subterm dag $G=(V,A,\lambda)$ for $E$ and $F$.
  2. 2.
    We compute the relations $INC(E\cup F,E)$, $INC(E\cup F,F)$, $NUM(E\cup F,E)$, $NUM(E\cup F,F)$, $BSTEP\langle E\cup F,\emptyset\rangle$, the GTRSs $R\langle E\cup F,\emptyset\rangle$, $R\langle E,F\rangle$, $R\langle F,E\rangle$, the values of the attributes $equal_{E}$, $equal_{F}$ for every vertex in $C\langle E,F\rangle$, and the satellite data $counter_{E}$, and $counter_{F}$ for every rule in $R\langle E\cup F,\emptyset\rangle$.
  3. 3.
    We compute the relations $NUM(E\cup F,E)$ and $NUM(E\cup F,F)$.
  4. 4.
    We create the adjacency-list representation of the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$.
  5. 5.
    We compute the values of the attribute $keeps_{E}$.
  6. 6.
    We compute the values of the attribute $keeps_{F}$.)
1.{ Create a subterm dag $G=(V,A,\lambda)$ for $E$ and $F$. }
@@
By Proposition 2.17, compute the congruence closure $\rho\langle F,E\rangle$ of the relation $\tau\langle F,E\rangle=\{\,(u,v)\mid(\hat{u},\hat{v})\in F\,\}.$
@@
for each ${d}\in C\langle E\cup F,\emptyset\rangle$ do
@@
${d}.equal_{E}=\mathrm{\mathrm{false}};$
${d}.equal_{F}=\mathrm{\mathrm{false}};$
@@
${b}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
${c}:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x)$;
${d}:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x)$;
$INC(E\cup F,E):=INC(E\cup F,E)\cup\{\,({d},{b})\,\}$;
(store $INC(E\cup F,E)$ without duplicates in the red-black tree $RBT1$.)
$INC(E\cup F,F):=INC(E\cup F,F)\cup\{\,({d},{c})\,\}$;
(store $INC(E\cup F,F)$ without duplicates in the red-black tree $RBT2$.)
if |b|=|d||{b}|=|{d}|
$\bf then$} ${d}.equal_{E}:=\mathrm{\mathrm{true}};$
if |c|=|d||{c}|=|{d}|
$\bf then$} ${d}.equal_{F}:=\mathrm{\mathrm{true}};$
(By Proposition 2.17 we know the cardinalities of all $W/_{\Theta\langle E,F\rangle}$-classes, $W/_{\Theta\langle F,E\rangle}$-classes, and $W/_{\Theta\langle E\cup F,\emptyset\rangle}$-classes.)
@@
${b}_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
${c}_{i}:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x/i)$;
${d}_{i}:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x/i)$;
@@
if $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\not\in R\langle E\cup F,\emptyset\rangle$
@@
$R\langle E\cup F,\emptyset\rangle:=R\langle E\cup F,\emptyset\rangle\cup\{\,\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\,\}$;
(Store $R\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT5$.)
$\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}:=0$;
$\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{F}:=0$;
$BSTEP\langle E\cup F,\emptyset\rangle:=BSTEP\langle E\cup F,\emptyset\rangle\cup\{\,({d}_{1},{d}),\ldots,({d}_{m},{d})\,\}$;
(Store $BSTEP\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT6$)
@@
if $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\not\in R\langle E,F\rangle$
@@
$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\,\}$;
(Store $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT7$.)
$\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}:=\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}+1$;